OpenStudy (anonymous):

Write a polynomial function with the given roots: 1 , -2 and 3i Be sure to show your work when writing equation.

OpenStudy (anonymous):

@Agl202

OpenStudy (agl202):

@siearra

OpenStudy (agl202):

Sorry, not sure about this one, @zeyatic

OpenStudy (anonymous):

lol It's a little odd.

OpenStudy (anonymous):

Complex roots always come in "pairs" - there's never a "solitary" complex root if I can put it that way. If a polynomial admits one complex root it also admits its conjugate. Therefore, your polynomial should have a minimum of 4 roots, these being 1,-2,3i and -3i. Working on that, we may construct our polynomial based on the general form of polynomials: P=a*(x-root1)*(x-root2)*.... , so forth depending on how many roots you have. where a is just about any constant except for 0 (duh) - although it could be if you want to be arrogant and call 0 a polynome. P=a*(x-1)*(x+2)*(x-3i)*(x+3i) in our case. where a can be just about anything (let it be 1 in this case) P=(x-1)*(x+2)*(x^2+9) P=(x^2+x-2)*(x^2+9) You work the final form.

OpenStudy (agl202):

WoW great job, dude. I have learned something new today... Thanks for helpin him... ; )

OpenStudy (anonymous):

No problem lol polynomials are fascinating.

OpenStudy (agl202):

Lol : )

OpenStudy (anonymous):

can you plz do that last part

OpenStudy (anonymous):

The last part is just about opening the parenthesis. Take each term from the left parenthesis and multiply it with each term of the right parenthesis.